OpenStudy (calculusxy):

1 mile = 1760 yards The distance from the house to the store is 3 miles, using the ratio, find the distance, in yards, from the house to the store. Show your work.

OpenStudy (calculusxy):

@phi

OpenStudy (anonymous):

3*1760=x*1 5280 =x 5280 yards

OpenStudy (ddcamp):

We're given that there are 1760 yards in a mile, and that it is 3 miles from the house to the store. Since the distance between the house and the store is 3*1 = 3 miles, there will be three times as many yards between the house and the store as there are in 1 mile. Thus, there are 3*1760 = 5280 yards between the house and the store.

OpenStudy (calculusxy):

What is the total distance in yards to make 3 trips from the house to the store and back in one week?

OpenStudy (ddcamp):

A trip from the house to the store and back includes two trips (one from the house to the store, one from the store to the house). This round-trip is twice the distance between the house and the store, or 2*5280 = 10560 yards. If there are 3 of these trips, then the total distance traveled is 3*10560 = 31680 yards.